# STBA: Towards Evaluating the Robustness of DNNs for Query-Limited Black-box Scenario

### v1

@@ IV-E Performance on Physical Online Vision Models
To evaluate the adversarial examples’ effectiveness in real-world scenarios, we conduct experiments of attacking the online models on Google Cloud Vision, Tencent Cloud and Baidu AI Cloud, whose structures and parameters and datasets used for training in their platforms are entirely sightless for us. We use the ImageNet dataset to carry out our attack on the pre-trained ResNet-152 model and then we randomly choose an adversarial image and its clean counterpart to validate on three online image recognition platforms. We draw the returned top-5 results in Fig. 6, which shows that the attack has successfully changed the image’s top-5 prediction results to a large degree. Taking Baidu AI Cloud for instance, its top-5 outputs of the clean image like Building (77.67%), Solar Thermal Collector (61.18%), Solar Panel (44.82%), Tile Roof (28.49%), Street (7.09%). However, after attacking, its top-5 predictions changed to Street (38.97%), Modern Architecture (32.94%), Swimming Pool (25.50%), Tree (13.43%) and Resort Hotel (1.17%). Besides, for Tencent Cloud, the adversarial example even increases the output labels from 3 to 4. This phenomenon indicates that our attack method is also applicable to the real physical world.
@@ IV-F Ablation Study
Figure 9: Ablation study of attack performance and generated adversarial examples’ quality under different hyper-parameters. The critical hyper-parameters involved here are the number of samples $n_{sample}$, the learning rate $lr$ and the number of adjustments of flow field budget $\xi$ $adjust_{num}$. We show the results of attack success rate, average query number and image quality in each row, respectively. The lines with different colors represent the result on different datasets, the solid lines represent the attack success rate, and the dashed lines indicate the average query number (Avg.Q) and image quality, respectively.
@@
In this section, we extend the ablation studies to verify the attack performance (including attack success rate and query times) and the generated adversarial examples’ visual quality under different hyper-parameters settings, where the visual quality is measured by the Learned Perceptual Image Patch Similarity (LPIPS) metric as the previous works [26, 28, 75].
@@
we first investigate the influence of loss function balancing factor $\lambda$ on the performance of the proposed method. Specifically, we conduct a dissection experiment to systematically vary the balancing factor $\lambda$ within the loss function and observe its effects on attack success rate and the image quality of the generated adversarial examples. The results Fig. 8 show that the attack can maintain a success rate when the $\lambda$ is set to a small value, i.e., 0.1 to 5, and has an obvious drop when $\lambda$ is greater than 5. While the image quality decreases with the $\lambda$ increase. Therefore, to balance the attack performance and image quality of the generated images, we set the $\lambda=5$ in this paper according to these empirical results.
@@
In addition, in our method, other involved critical hyper-parameters are as follows: the number of samples $n_{sample}$ to estimate the gradient, the learning rate $lr$ and the number of adjustments of flow field budget $\xi$, $adjust_{num}$. First of all, we set $\sigma=2*lr$ and do the experiments under a predefined hyper-parameters set; we illustrate the results in Fig. 9. Specifically, we set the $n_{sample}\in\{10,20,30,40\}$, $lr\in\{0.25,0.1,0.08,0.04,0.025,0.01\}$ and $adjust_{num}\in\{10,20,30,40\}$ and extent attacks on the ResNet (ResNet-56 for CIFAR and STL-10 dataset, and ResNet-152 for ImageNet dataset, respectively.) under the max query number is set to 1000. As the results show, for $n_{sample}$, we set $n_{sample}=10$ to get the highest attack success rate and lowest average query number and set $n_{sample}=30$ to get the best image quality, but the attack success rate will be faded for most victim datasets. For $lr$, we set $lr=0.1$ to balance the attack success rate, average query number, and image quality. For $adjust_{num}$, STBA achieves the best attack performance and image quality when $adjust_{num}=30$. Therefore, in our experiments, to obtain the adversarial images with high attack success and image quality, we set the most essential hyper-parameters as $n_{sample}=10$, $lr=0.1$ and $adjust_{num}=20$, respectively.
@@ V Conclusions and Limitations
To evaluate the vulnerability and robustness of existing normally trained and robustly trained DNN models under query-limited black-box situations, in this paper, we present a novel non-noise additional method, called STBA, to build adversarial examples. The proposed STBA combines performing the spatial transformation and gradient estimation to search for the best flow field to synthesize adversarial examples. Empirically, the adversarial examples generated by STBA are imperceptible to the human eyes and significantly increase query efficiency; it only needs dozens and hundreds of queries to obtain a practical adversarial example both in attacking normally trained and robustly trained models on four different benchmark datasets in most cases. Extensive experiments show that the proposed method is superior to the existing methods in terms of attack success rate, query times and imperceptibility. Even on robust models, it can obtain 93.27 $\sim$ 100% attack success rate on benchmark datasets. Benefitting from generating adversarial examples without noise-adding, the proposed STBA provides a new efficient way to evaluate the vulnerability of existing classifiers and further enhance their robustness performance using techniques like fine-tuning or adversarial training. Furthermore, the proposed approach can be a reliable tool for building new robust DNN models.
@@ Acknowledgments
This work is supported in part by the National Natural Science Foundation of China under Grant 62162067 and 62101480, and in part by the Yunnan Province expert workstations under Grant 202305AF150078.
@@
[Uncaptioned image] Renyang Liu received the BE degree in Computer Science from the Northwest Normal University in 2017. He is a current Ph.D. candidate in the School of Information Science and Engineering at Yunnan University, Kunming, China. Now, he is studying at the School of Computer Science and Engineering at Nanyang Technological University as a Joint-training Ph.D. student. His current research interest includes deep learning, AI security and privacy preserving.
[Uncaptioned image] Kwok-Yan Lam (Senior Member, IEEE) received his B.Sc. degree (1st Class Hons.) from University of London, in 1987, and Ph.D. degree from University of Cambridge, in 1990. He is the Associate Vice President (Strategy and Partnerships) and Professor in the School of Computer Science and Engineering at the Nanyang Technological University, Singapore. He is currently also the Director of the Strategic Centre for Research in Privacy-Preserving Technologies and Systems (SCRiPTS). From August 2020, he is on part-time secondment to the INTERPOL as a Consultant at Cyber and New Technology Innovation. Prior to joining NTU, he has been a Professor of the Tsinghua University, PR China (2002–2010) and a faculty member of the National University of Singapore and the University of London since 1990. He was a Visiting Scientist at the Isaac Newton Institute, Cambridge University, and a Visiting Professor at the European Institute for Systems Security. In 1998, he received the Singapore Foundation Award from the Japanese Chamber of Commerce and Industry in recognition of his research and development achievement in information security in Singapore. His research interests include Distributed Systems, Intelligent Systems, IoT Security, Distributed Protocols for Blockchain, Homeland Security and Cybersecurity.
@@
[Uncaptioned image] Jun Zhao (Member, IEEE) is currently an Assistant Professor in the School of Computer Science and Engineering (SCSE) at Nanyang Technological University (NTU), Singapore. He received a Ph.D. degree in Electrical and Computer Engineering from Carnegie Mellon University (CMU), Pittsburgh, PA, USA, in May 2015, and a bachelor’s degree in Information Engineering from Shanghai Jiao Tong University, China, in June 2010. One of his papers was a finalist for the best student paper award in IEEE International Symposium on Information Theory (ISIT) 2014. His research interests include A.I. and data science, security and privacy, control and learning in communications and networks.

### v2

@@ IV-E Performance on Physical Online Vision Models
To evaluate the adversarial examples’ effectiveness in real-world scenarios, we conduct experiments of attacking the online models on Google Cloud Vision, Tencent Cloud and Baidu AI Cloud, whose structures and parameters and datasets used for training in their platforms are entirely sightless for us. We use the ImageNet dataset to carry out our attack on the pre-trained ResNet-152 model and then we randomly choose an adversarial image and its clean counterpart to validate on three online image recognition platforms. We draw the returned top-5 results in Fig. 6, which shows that the attack has successfully changed the image’s top-5 prediction results to a large degree  [^12] We call these APIs on March 24, 2024.. Taking Baidu AI Cloud for instance, its top-5 outputs of the clean image like Building (77.67%), Solar Thermal Collector (61.18%), Solar Panel (44.82%), Tile Roof (28.49%), Street (7.09%). However, after attacking, its top-5 predictions changed to Street (38.97%), Modern Architecture (32.94%), Swimming Pool (25.50%), Tree (13.43%) and Resort Hotel (1.17%). Besides, for Tencent Cloud, the adversarial example even increases the output labels from 3 to 4. This phenomenon indicates that our attack method is also applicable to the real physical world.
@@ IV-F Ablation Study
TABLE V: The attack performance, query efficiency, and visual quality of applying the spatial transform to High- $vs.$ Low-frequency part.
Dataset Frequency ASR Avg.Q Med.Q LPIPS DISTS
CIAFAR-10 High 99.40 142.47 111.00 0.02 0.05
Low 31.40 387.19 342.00 0.05 0.07
CIFAR-100 High 99.60 132.58 89.00 0.02 0.05
Low 48.10 363.68 298.00 0.06 0.08
STL-10 High 99.10 241.65 190.00 0.02 0.06
Low 43.40 385.62 337.00 0.05 0.10
ImageNet High 99.50 205.34 169.00 0.02 0.06
Low 27.70 293.26 169.00 0.06 0.09
@@
Figure 9: Ablation study of attack performance and generated adversarial examples’ visual quality under different hyper-parameters. The critical hyper-parameters involved here are the number of samples $n_{sample}$, the learning rate $lr$ and the number of adjustments of flow field budget $\xi$ $adjust_{num}$. We show the results of attack success rate, average query number and image quality in each row, respectively. The lines with different colors represent the result on different datasets, the solid lines represent the attack success rate, and the dashed lines indicate the average query number (Avg.Q) and image quality, respectively.
@@
In this section, we extend the ablation studies to verify the attack performance (including attack success rate and query times) and the generated adversarial examples’ visual quality under different hyper-parameter settings.
We first investigate the effectiveness of the proposed strategy in terms of applying the spatial transform to the high-frequency part to improve the attack performance and the visual quality of generated adversarial examples. We list the results in Table V, which verified our claim and shows that applying spatial transform to the high-frequency part not only gets a good attack performance but also improves the query efficiency and image quality. The test images are randomly selected and correctly classified 1000 images for CIFAR-10, CIFAR-100, and STL-10, while the ImageNet (NIPS) are all correctly classified images. Here, the maximal query budget is set to 1000, and the victim model is VGG-19.
@@
Besides, we also investigate the influence of loss function balancing factor $\lambda$ on the performance of the proposed method. Specifically, we conduct a dissection experiment to systematically vary the balancing factor $\lambda$ within the loss function and observe its effects on attack success rate and the image quality of the generated adversarial examples. The results Fig. 8 show that the attack can maintain a success rate when the $\lambda$ is set to a small value, i.e., 0.1 to 5, and has an obvious drop when $\lambda$ is greater than 5. While the image quality decreases with the $\lambda$ increase. Therefore, to balance the attack performance and image quality of the generated images, we set the $\lambda=5$ in this paper according to these empirical results.
@@
Finally, in our method, other involved critical hyper-parameters are as follows: the number of samples $n_{sample}$ to estimate the gradient, the learning rate $lr$ and the number of adjustments of flow field budget $\xi$, $adjust_{num}$. First of all, we set $\sigma=2*lr$ and do the experiments under a predefined hyper-parameters set; we illustrate the results in Fig. 9. Specifically, we set the $n_{sample}\in\{10,20,30,40\}$, $lr\in\{0.25,0.1,0.08,0.04,0.025,0.01\}$ and $adjust_{num}\in\{10,20,30,40\}$ and extent attacks on the ResNet (ResNet-56 for CIFAR and STL-10 dataset, and ResNet-152 for ImageNet dataset, respectively.) under the max query number is set to 1000. As the results show, for $n_{sample}$, we set $n_{sample}=10$ to get the highest attack success rate and lowest average query number and set $n_{sample}=30$ to get the best image quality, but the attack success rate will be faded for most victim datasets. For $lr$, we set $lr=0.1$ to balance the attack success rate, average query number, and image quality. For $adjust_{num}$, STBA achieves the best attack performance and image quality when $adjust_{num}=30$. Therefore, in our experiments, to obtain the adversarial images with high attack success and image quality, we set the most essential hyper-parameters as $n_{sample}=10$, $lr=0.1$ and $adjust_{num}=20$, respectively.
@@ V Conclusions and Limitations
To evaluate the vulnerability and robustness of existing normally trained and robustly trained DNN models under query-limited black-box situations, in this paper, we present a novel non-noise additional method, called STBA, to build adversarial examples. The proposed STBA combines performing the spatial transformation to the high-frequency part of the image and gradient estimation to search for the best flow field to synthesize adversarial examples. Empirically, the adversarial examples generated by STBA are imperceptible to the human eyes and significantly increase query efficiency; it only needs dozens and hundreds of queries to obtain a practical adversarial example in attacking normally trained and robustly trained models, respectively. Extensive experiments show that the proposed method is superior to the existing methods in terms of attack success rate, query times and imperceptibility. Even on robust models, it can obtain 93.27 $\sim$ 100% attack success rate on benchmark datasets. Benefitting from generating adversarial examples without noise-adding, the proposed STBA provides a new efficient way to evaluate the vulnerability of existing classifiers and further enhance their robustness performance using techniques like fine-tuning or adversarial training.
@@ Acknowledgments
Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of the National Research Foundation, Singapore and Infocomm Media Development Authority, Yunnan Province expert workstations, Yunnan Fundamental Research Projects.
@@
[Uncaptioned image] Renyang Liu (Graduate Student Member, IEEE) earned his B.E. in Computer Science from Northwest Normal University in 2017 and his Ph.D. from Yunnan University in 2024. He served as a Joint-training Ph.D. student at Nanyang Technological University’s College of Computing and Data Science from 2022 to 2023. Currently, he is a Research Fellow at the Institute of Data Science, National University of Singapore. His research primarily focuses on security issues of multimodal models, with interests spanning AI security, machine unlearning, data privacy, and computer vision.
[Uncaptioned image] Kwok-Yan Lam (Senior Member, IEEE) received his B.Sc. degree (1st Class Hons.) from University of London, in 1987, and Ph.D. degree from University of Cambridge, in 1990. He is the Associate Vice President (Strategy and Partnerships) and Professor in the College of Computing and Data Science at the Nanyang Technological University, Singapore. He is currently also the Director of the Strategic Centre for Research in Privacy-Preserving Technologies and Systems (SCRiPTS). From August 2020, he is on part-time secondment to the INTERPOL as a Consultant at Cyber and New Technology Innovation. Prior to joining NTU, he has been a Professor of the Tsinghua University, PR China (2002–2010) and a faculty member of the National University of Singapore and the University of London since 1990. He was a Visiting Scientist at the Isaac Newton Institute, Cambridge University, and a Visiting Professor at the European Institute for Systems Security. In 1998, he received the Singapore Foundation Award from the Japanese Chamber of Commerce and Industry in recognition of his research and development achievement in information security in Singapore. His research interests include Distributed Systems, Intelligent Systems, IoT Security, Distributed Protocols for Blockchain, Homeland Security and Cybersecurity.
@@
[Uncaptioned image] Jun Zhao (Member, IEEE) is currently an Assistant Professor in the College of Computing and Data Science at Nanyang Technological University (NTU), Singapore. He received a Ph.D. degree in Electrical and Computer Engineering from Carnegie Mellon University (CMU), Pittsburgh, PA, USA, in May 2015, and a bachelor’s degree in Information Engineering from Shanghai Jiao Tong University, China, in June 2010. One of his papers was a finalist for the best student paper award in IEEE International Symposium on Information Theory (ISIT) 2014. His research interests include A.I. and data science, security and privacy, control and learning in communications and networks.